A projectile lands 250 m from the base of a vertical cliff 150 m high. How fast was it launched?


1
Expert's answer
2020-10-12T07:46:39-0400

The distance travelled by the object in the vertical direction is:


"h = \dfrac{gt^2}{2}"

If the height is "h = 150m" and "g = 9.81m\/s^2" is the gravitational acceleration, then the falling time will be:


"t_{fall} = \sqrt{\dfrac{2h}{g}}"

Object's horizontal speed has not been changed with time. Thus, the distance in the horizontal direction travelled in time of falling will be:


"d = v_0t_{fall} = v_0\sqrt{\dfrac{2h}{g}}"


where "v_0" is the initial velocity. If "d = 250m", then the initial velocity is:


"v_0 = d\sqrt{\dfrac{g}{2h}} = 250\sqrt{\dfrac{9.81}{2\cdot 150}}\approx 45.2m/s"

Answer. 45.2 m/s.